JosephGunnels:

h(x) = x2 + 1 k(x) = x – 2 (h + k)(2) =

supie:

Ok so, \[ h(x) = x²+1 \ and\ k(x) = x-2\] \[h(x)+k(x) = (x²+1)+(x-2)\] \[h(x)+k(x) = x²+x-1\] \[To\ get (h+k)(2), we\ will\ substitute x=2 into\ the\ resulting\ function\ to\ give\] \[(h+k)(2) = 2²+2-1\] \[(h+k)(2) = 6-1\] \[(h+k)(2) =?\]

xxDeppressionxx:

(h+k) =5 (h-k) (3)=9 3h(2) +2k(3)=17

Timmyspu:

@supie is right so after that then what do you do

Prankster233Mya:

oki then?

Timmyspu:

@xxDeppressionxx u can not give a dirrect awnser u have to let him/her figure it out

xxDeppressionxx:

i didnt

xxDeppressionxx:

i explained it

supie:

\(\color{#0cbb34}{\text{Originally Posted by}}\) @xxDeppressionxx i explained it \(\color{#0cbb34}{\text{End of Quote}}\) Yeah he explained it.

JosephGunnels:

would it be 5 and I am a him

Timmyspu:

ok woops my bad

JosephGunnels:

its ok

supie:

\(\color{#0cbb34}{\text{Originally Posted by}}\) @JosephGunnels would it be 5 and I am a him \(\color{#0cbb34}{\text{End of Quote}}\) Yeah good job.

MrMudd183:

good one supie, nice explanation of the work and step-by-step process

xxDeppressionxx:

ur good but next time dont forget that when someone explains the answer they got its not giving a direct answer

supie:

\(\color{#0cbb34}{\text{Originally Posted by}}\) @MrMudd183 good one supie, nice explanation of the work and step-by-step process \(\color{#0cbb34}{\text{End of Quote}}\) Thanks Mudd!

MrMudd183:

You're welcome :)

JosephGunnels:

thank u i got another question to post hope yall can help